# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v4

@@ Abstract
This work presents two significant contributions from the perspectives of QRNG manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
@@ 1 Introduction
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming [1, 2], testing principles of physics and chemistry [3], as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages [4, 5], extending to commercial uses such as slot machines and lottery games [6, 7]. In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source [8, 9], which may be classical or quantum.
@@
The randomness of a classical true random number generator (CTRNG) depends on the unpredictability of the physical process that generates the entropy. This unpredictability arises from our limited knowledge of the system’s behavior [9, 10]. However, if future research resolves these limitations, the process could become predictable and would no longer function as a true random number generator. In contrast, a quantum true random number generator (QTRNG, or more commonly, QRNG) exploits the properties of quantum particles and the principles of quantum mechanics [11] to produce random numbers. According to quantum theory, these numbers are fundamentally random, making them inherently unpredictable. Thus it would remain unpredictable, irrespective of any technological advancement.
@@
Different technologies, including photon arrival time [12, 13, 14, 15, 16, 17, 18, 19], tunneling [20, 21], fluctuation [22, 23, 24, 25, 26, 27], phase noise [28, 29, 30, 31, 32, 33, 34], etc. have been proposed to produce QRNGs [35, 10]. Several manufacturers such as ID Quantique [36], Qutools [37], QNuLabs [38], QuintessenceLabs [39], Mars Innovation [40], Quantum Computing Inc. [41], Quside [42], QuantumCTek [43] are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
@@
Since no black-box randomness-testing algorithm can distinguish between a PRNG and a QRNG, therefore the first step in validating the quantumness [^1] Sub-Poissonian distribution ($mean>variance$) of photon statistics (more details in Section 2). of a random number generator (RNG) is to collect information regarding intermediate attenuator and detector parameters and photon statistics. Then it may be possible to devise a strategy to decide whether the device performs in the classical or quantum regime.
@@
In most of the analyses [44, 45, 46, 47] of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter [48]. For example, Alléaume et al. [49] discussed photon statistics of photon sources using the Mandel parameter given by [48, Eq. (11a)]. They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources.
@@
Although the above strategy is analytically correct, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same (even under approximation) as the mean and the variance of the underlying distribution [50]. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
@@
Another approach to test the quantumness of light is to observe photon antibunching [51] experimentally. However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons [52, 53]. We explain this in more detail in Appendix B. This approach also consists of equality criteria similar to the above photon statistics case. Therefore this is also not enough to decide photon antibunching.
@@
On the other hand, a single-photon source ensures the quantumness of a QRNG [51, 53]. However, generating single photons are experimentally challenging task [54, 55]. Various experimental methods [56, 57, 58, 59, 60, 61, 62, 63] have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential [15, 14, 16] or uniform [17] distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Nie et al. [17] shows that, for a photon arrival-time-based QRNG, the uniformity and hence the min-entropy depends on device parameters such as expected photon number within a fixed time interval as well as on the detection efficiency. Therefore, it is important to explore photon statistics in practical scenarios in more detail.
@@
1.1 Motivation
Given that both PRNGs and QRNGs produce sequences of 0’s and 1’s, merely analyzing these sequences is not sufficient to definitively differentiate between the two [64]. Statistical tests like NIST [65], Dieharder [66], AIS-$31$ [67], ENT [68] etc. are available to test the randomness of an RNG. However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
@@ 1.1 Motivation
Also, there are several schemes of device-independent QRNG (DI-QRNG) [69, 70, 71, 72, 73, 74, 75] and certified randomness [69, 76, 77, 71, 78, 79, 80, 19, 81, 82, 83] based on Bell violation, non-separability, etc. [71, 84, 85, 86]. Though they do not rely on trusted device assumptions and do not require any post-processing, the required resources (that is, entangled qubits) are too costly. More recent works [87, 88] attempt locally certified randomness without entanglement, but they require random seeds and some kind of bootstrapping that may be considered analogous to classical post-processing.
On the other hand, the optical trusted device QRNG [10] relies on the assumption that the devices used for extracting randomness and the source are well-characterized and no one can change any configuration. This means that no third party can access the QRNG or influence its results. In contrast, single-photon-based QRNGs can still be verified using photon-count statistics [52, 53, 89, 90] improving the security. A detailed comparison between trusted QRNG and DI-QRNG is given in Appendix A. In this work, we focus on the QRNG of the optical trusted device type along with a two-fold statistical method to verify its quantumness.
Side-channel attacks [91, 92, 93] can be performed in the post-processing step to gather information about the random number. Therefore, to reduce the compression amount of raw data and to prevent the data from side-channel attacks, the amount of post-processing should be minimal to ensure the maximum entropy in raw data with high speed and reduce the possibility of side-channel attacks on post-processing. To achieve this, the raw output number (before post-processing) generated by QRNG should be as close to uniform random as possible. Nie et al. [17] provided a lower bound for the min-entropy of the raw data in the presence of device errors. Depending on this min-entropy, a suitable post-processing algorithm can be chosen to retain the entropy from the quantum source. As all the practical devices are prone to error, this raises an important question:
@@
1.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level.
@@ 1.2 Our Contributions
Finally, we simulated raw data depending on photon statistics we have derived and performed statistical tests like NIST [65], Dieharder [66], AIS-$31$ [67] and ENT [68] on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. [17] performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
This paper is structured as follows. In Section 2 we discuss the quantumness of a single-photon source. In Section 3 and 4 we explore single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. We also discuss noisy beam-splitters and detectors in the case of single-photon-based QRNG in Section 3. Section 5 reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section 6 to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. [17]. Section 7 answers Q3 by separately analyzing the effects of the parameters mentioned in that question. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section 6 are given in Section 8. Finally, in Section 9 we conclude our work.
2 Quantumness of a Single-photon Source
In this section, we describe what we mean by quantumness in the case of a single-photon source. We first discuss photon statistics and then revisit the existing procedure to distinguish quantum light from classical light using such photon statistics. Here, we validate the quantumness of a photon source by its statistical closeness to an ideal single-photon source. Another way of defining and validating quantumness is through Bell test, quantum discord, non-separability, etc. [71, 84, 85, 86] for QRNGs that are based on entangled photon source. As our point of interest is the single-photon source which does not have any entanglement, we are not concerned with these tests in our current work.
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols [4, 5]. In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, photon-counting statistics play an important role in determining whether or not light falls in the quantum regime.
The distribution of the single-photon-counting statistics can be divided into three classes [53]:
@@ 2 Quantumness of a Single-photon Source
Classical wave theory of light follows Poissonian and super-Poissonian statistics [53]. On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature [89, 90]. A detailed discussion is given in Appendix B.
@@
Thermal light follows Bose–Einstein distribution [53], given by
@@
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution [94] given by
@@
Antibunching property can also be used to distinguish quantum light from classical light. Detailed discussion about this is included in Appendix B.
3 QRNG based on Single-photon Source
An ideal QRNG, based on single photons, uses $/50$ 50 beam splitter [95], where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.

### v5

@@ Abstract
This work presents two significant contributions from the perspectives of quantum random number generator (QRNG) manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
@@ I Introduction
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming RevModPhys.87.1067 ; MonteCarlo , testing principles of physics and chemistry PhysicsPhysiqueFizika.1.195 , as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett ; Crypto , extending to commercial uses such as slot machines and lottery games konstantinou2004electronic ; yamanaka2017enumeration . In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source Schindler2009 ; Stipčević2014 , which may be classical or quantum.
@@
The randomness of a classical true random number generator (CTRNG) depends on the unpredictability of the physical process that generates the entropy. This unpredictability arises from our limited knowledge of the system’s behavior Stipčević2014 ; mannalatha2023comprehensive . However, if future research resolves these limitations, the process could become predictable and would no longer function as a true random number generator. In contrast, a quantum true random number generator (QTRNG, or more commonly, QRNG) exploits the properties of quantum particles and the principles of quantum mechanics ma2016quantum to produce random numbers. According to quantum theory, these numbers are fundamentally random, making them inherently unpredictable. Thus it would remain unpredictable, irrespective of any technological advancement.
@@
Different technologies, including photon arrival time 10.1063/1.2720728 ; 10.1063/1.2961000 ; WayneRNG ; WayneQRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 ; 7270265 ; zhang2021simple , tunneling bernardo2017extracting ; 10.1063/5.0055955 , fluctuation gabriel2010generator ; PhysRevApplied.3.054004 ; Guo:19 ; ZQVR ; 10.1063/1.5078547 ; gehring2021homodyne , phase noise Qi:10 ; PhysRevE.81.051137 ; PhysRevA.81.063814 ; Xu:12 ; 10.1063/1.4922417 ; Yang:16 ; app10072431 , etc. have been proposed to produce QRNGs RevModPhys.89.015004 ; mannalatha2023comprehensive . Several manufacturers such as ID Quantique quantis , Qutools quRNG , QNuLabs TROPOS , QuintessenceLabs qStream , Mars Innovation quantoss , Quantum Computing Inc. uQRNG , Quside quside , QuantumCTek QCTek are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
@@
Since no black-box randomness-testing algorithm can distinguish between a PRNG and a QRNG, therefore the first step in validating the quantumness [^1] Sub-Poissonian distribution ($mean>variance$) of photon statistics (more details in Section II). of a random number generator (RNG) is to collect information regarding intermediate attenuator and detector parameters and photon statistics. Then it may be possible to devise a strategy to decide whether the device performs in the classical or quantum regime.
@@
In most of the analyses Jones:23 ; Perina:24 ; Perina24 ; PhysRevA.109.043505 of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter Mandel:79 . For example, Alléaume et al. Alléaume_2004 discussed photon statistics of photon sources using the Mandel parameter given by (Mandel:79, , Eq. (11a)). They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources.
@@
Although the above strategy is analytically correct, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same (even under approximation) as the mean and the variance of the underlying distribution BERG2005865 . Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
@@
Another approach to test the quantumness of light is to observe photon antibunching RevModPhys.54.1061 experimentally. However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons PhysRevA.41.475 ; fox2006quantum . We explain this in more detail inB. This approach also consists of equality criteria similar to the above photon statistics case. Therefore this is also not enough to decide photon antibunching.
@@
On the other hand, a single-photon source ensures the quantumness of a QRNG RevModPhys.54.1061 ; fox2006quantum . However, generating single photons are experimentally challenging task PhysRevLett.72.210 ; kim1999single . Various experimental methods PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential WayneQRNG ; WayneRNG ; 10.1063/1.3578456 or uniform 10.1063/1.4863224 distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Nie et al. 10.1063/1.4863224 shows that, for a photon arrival-time-based QRNG, the uniformity and hence the min-entropy depends on device parameters such as expected photon number within a fixed time interval as well as on the detection efficiency. Therefore, it is important to explore photon statistics in practical scenarios in more detail.
@@
I.1 Motivation
Given that both PRNGs and QRNGs produce sequences of 0’s and 1’s, merely analyzing these sequences is not sufficient to definitively differentiate between the two PhysRevA.109.022243 . Statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal , ENT ENT etc. are available to test the randomness of an RNG. However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
@@ I.1 Motivation
Also, there are several schemes of device-independent QRNG (DI-QRNG) pironio2010random ; 10.1145/2885493 ; acin2016certified ; bierhorst2018experimentally ; liu2018device ; PhysRevLett.120.010503 ; 9153851 and certified randomness pironio2010random ; 6326426 ; Passaro_2015 ; acin2016certified ; Nguyen:18 ; 10.1093/oso/9780198788416.003.0008 ; PhysRevX.10.041048 ; zhang2021simple ; Leone:23 ; PhysRevLett.133.020802 ; PRXQuantum.5.020348 based on Bell violation, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . Though they do not rely on trusted device assumptions and do not require any post-processing, the required resources (that is, entangled qubits) are too costly. More recent works Mironowicz_2016 ; 10.1145/3564246.3585145 attempt locally certified randomness without entanglement, but they require random seeds and some kind of bootstrapping that may be considered analogous to classical post-processing.
On the other hand, the optical trusted device QRNG mannalatha2023comprehensive relies on the assumption that the devices used for extracting randomness and the source are well-characterized and no one can change any configuration. This means that no third party can access the QRNG or influence its results. In contrast, single-photon-based QRNGs can still be verified using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 improving the security. A detailed comparison between trusted QRNG and DI-QRNG is given in A. In this work, we focus on the QRNG of the optical trusted device type along with a two-fold statistical method to verify its quantumness.
Side-channel attacks 8924447 ; 8942152 ; chowdhury2021physical can be performed in the post-processing step to gather information about the random number. Therefore, to reduce the compression amount of raw data and to prevent the data from side-channel attacks, the amount of post-processing should be minimal to ensure the maximum entropy in raw data with high speed and reduce the possibility of side-channel attacks on post-processing. To achieve this, the raw output number (before post-processing) generated by QRNG should be as close to uniform random as possible. Nie et al. 10.1063/1.4863224 provided a lower bound for the min-entropy of the raw data in the presence of device errors. Depending on this min-entropy, a suitable post-processing algorithm can be chosen to retain the entropy from the quantum source. As all the practical devices are prone to error, this raises an important question:
@@
I.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level. Several photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years that can be used to perform these tests.
@@ I.2 Our Contributions
Finally, we simulated raw data depending on photon statistics we have derived and performed statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal and ENT ENT on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. 10.1063/1.4863224 performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
This paper is structured as follows. In Section II we discuss the quantumness of a single-photon source. In Section III and IV we explore single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. We also discuss noisy beam-splitters and detectors in the case of single-photon-based QRNG in Section III. Section V reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section VI to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. 10.1063/1.4863224 . Section VII answers Q3 by separately analyzing the effects of the parameters mentioned in that question. The answer to Q3 by combining the randomness with random number generation rate and entropy for maximum rate is presented in Section VIII. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section VI are given in Section IX. Finally, in Section X we conclude our work.
II Quantumness of a Single-photon Source
In this section, we describe what we mean by quantumness in the case of a single-photon source. We first discuss photon statistics and then revisit the existing procedure to distinguish quantum light from classical light using such photon statistics. Here, we validate the quantumness of a photon source by its statistical closeness to an ideal single-photon source. Another way of defining and validating quantumness is through Bell test, quantum discord, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 for QRNGs that are based on entangled photon source. As our point of interest is the single-photon source which does not have any entanglement, we are not concerned with these tests in our current work.
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols Bennett ; Crypto . In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, photon-counting statistics play an important role in determining whether or not light falls in the quantum regime.
The distribution of the single-photon-counting statistics can be divided into three classes fox2006quantum :
@@ II Quantumness of a Single-photon Source
Classical wave theory of light follows Poissonian and super-Poissonian statistics fox2006quantum . On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature RevModPhys.68.127 ; qute.202100062 . A detailed discussion is given in B.
@@
Thermal light follows Bose–Einstein distribution fox2006quantum , given by
@@
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution PhysRev.145.1041 given by
@@
Antibunching property can also be used to distinguish quantum light from classical light. Detailed discussion about this is included in B.
III QRNG based on Single-photon Source
An ideal QRNG, based on single photons, uses $/50$ 50 beam splitter jennewein2000fast , where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.